OpenStudy (anonymous):

1/4log3 81+ 1/2log3 9- log3 27

OpenStudy (anonymous):

\[\frac{1}{4}\log_3(81)=\log_3(\sqrt[4]{81})=\log_3(3)=1\]

OpenStudy (anonymous):

similarly \[\frac{1}{2}\log_3(9)=\log_3(\sqrt{9})=\log_3(3)=1\]

OpenStudy (anonymous):

and finally \[\log_3(27)=\log_3(3^3)=3\]

OpenStudy (anonymous):

so "final answer" is \[1+1-3=-1\]

OpenStudy (anonymous):

wow thank you so much!!

OpenStudy (anonymous):

yw